13. The half ife of cesium-137 is 30 years. Suppose we have a 25g sample. a. (1 that models the mass remaining after t years l point) Find a function m)- m,2 13a b. (2 points) Find a function me)me- that models the mass remaining after t years In 2 (recall that r I), Do Not round. 13b: 2 points) How much of the sample will remain after 95 years? Round to the nearest ten-thousandth C. ( 13c:,
Answer #1

Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

a)

m0=25g

Since half life is 30

So, m0/2=m0*2^(-15/h)

1=15/h

h=15

So,

m=25*2^(-t/15)

b)

m0=25g

Since half life is 30

So, m0/2=m0*exp(-r*15)

log(2)=r*15

r=log(2)/15

So,

m=25*exp(-t*log(2)/15)

c)

m=25*exp(-t*log(2)/15)

t=95

So,

m=0.31003 g remaining

Kindly revert for any queries

Thanks.